OpenStudy (anonymous):

heeellpppp pleasseeeee ! 3(x+2)^2-10(x+6)-8=0 expand and collect like terms in the original equation , and solve the resulting equation for x??

OpenStudy (anonymous):

3(x+2)^2-10(x+6)-8 3(x^2+4x+4)-10x-60-8 3x^2+12x+12-10x-68 3x^2+2x-56 (3x+14)(x-4) x=-14/3,4

OpenStudy (anonymous):

Sorry for how long it took, I kept making mistakes. My numbers kept running together bee n looking at computer for too long. lol

OpenStudy (anonymous):

thanks so much and its ok i just needed the help i couldnt figure it out.

OpenStudy (anonymous):

-14 over 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

something wrong

OpenStudy (anonymous):

hmmm...im not sure where is.

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

I am not sure what is wrong. What I copied and solved is correct. I checked multiple times. Did you maybe copy it wrong?

OpenStudy (anonymous):

yeah i did sorry its supposed to be 3(x+6)^2-10(x+6)-8=0

OpenStudy (anonymous):

can you factor out (x+6)?

OpenStudy (anonymous):

ahh i see the problem, what you origianlly wrote had a 2 where 6 is

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

3(x+6)^2-10(x+6)-8 3(x^2+12x+36)-10x-60-8 3x^2+36x+108-10x-68 3x^+26x+40 (3x+20)(x+2) x=-20/3,-2

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

haha it is not a problem